OpenStudy (anonymous):

5. 3.75x + 3.7 = 1.7 + 1.75x (1 point) 10 –1 1 .1

OpenStudy (anonymous):

3.75x+3.7=1.7+1.75x subtract 1.75x from both sides 2x+3.7=1.7 subtract 3.7 from both sides 2x=-2 divide by two x=-1